3¾ ÷ ¾ is equal to:
(a) 3 (b) 4 c) 5 (d) 45/16
3¾ ÷ ¾ is equal to:
(a) 3 (b) 4 c) 5 (d) 45/16

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Option (c) is correct here
First we have to convert the mixed fraction into improper fraction 33⁄4 = 15/4
Then, 15/4 ÷ 3⁄4

= (15/4)/ (3⁄4)
= (15/4) × (4/3)
= (15 × 4)/ (4 ×3)
= (5 × 1)/ (1 ×1)
=5


Thanks